### v2

# The number of edges in graphs with bounded clique number and circumference

Abstract
Let ${\cal H}$ be a family of graphs. The Turán number ${\rm ex}(n,{\cal H})$ is the maximum possible number of edges in an $n$-vertex graph which does not contain any member of ${\cal H}$ as a subgraph. As a common generalization of Turán’s theorem and Erdős-Gallai theorem on the Turán number of matchings, Alon and Frankl determined ${\rm ex}(n,{\cal H})$ for =${\cal H}=\{K_{r},M_{k}\}$, where $M_{k}$ is a matching of size $k$. Replacing $M_{k}$ by $P_{k}$, Katona and Xiao obtained the Turán number of =${\cal H}=\{K_{r},P_{k}\}$ for $r\leq\lfloor k/2\rfloor$ and sufficiently large $n$. In addition, they proposed a conjecture for the case of $r\geq\lfloor k/2\rfloor+1$ and sufficiently large $n$. Motivated by the fact that the result for ${\rm ex}(n,P_{k})$ can be deduced from the one for ${\rm ex}(n,{\cal C}_{\geq k})$, we investigate the Turán number of =${\cal H}=\{K_{r},{\cal C}_{\geq k}\}$ in this paper. Namely, for such ${\cal H}$, we are able to show the value of ${\rm ex}(n,{\cal H})$ for $r\geq\lfloor(k-1)/2\rfloor+2$ and all $n$. As an application of this result, we confirm Katona and Xiao’s conjecture in a stronger form. For $r\leq\lfloor(k-1)/2\rfloor+1$, we manage to show the value of ${\rm ex}(n,{\cal H})$ for sufficiently large $n$.
Keywords:  Turán number; circumference; cycle
1 Introduction
The study of Turán number of graphs is one of the central topics in extremal graph theory. For a family of graph ${\cal H}$, a graph is ${\cal H}$-free if it does not contain a member of ${\cal H}$ as a subgraph. The Turán number ${\rm ex}(n,{\cal H})$ is the maximum possible number of edges in an $n$-vertex ${\cal H}$-free graph. If ${\cal H}$ contains one graph $H$, then we write ${\rm ex}(n,H)$ instead of ${\rm ex}(n,{\cal H})$. Mantel Theorem states that ${\rm ex}(n,K_{3})=\lfloor n^{2}/4\rfloor$ and the only extremal graph is the balanced bipartite graph. For cliques, the famous Turán Theorem tells us that ${\rm ex}(n,K_{p+1})=e(T(n,p))$ and the only extremal graph is $T(n,p)$, where $T(n,p)$ is the balanced complete $p$-partite graph with $n$ vertices. The number of edges in $T(n,p)$ is denoted by $t(n,p)$. Turán’s Theorem is viewed as the origin of extremal graph theory. For the matching $M_{k}$ of size $k$, Erdős and Gallai [7] determined the value of ${\rm ex}(n,M_{k})$. Let $P_{k}$ be the path with $k$ vertices and ${\cal C}_{\geq k}$ be the set of cycles of length at least $k$. In the same paper, Erdős and Gallai proved the following results.
Theorem 1 (Erdős-Gallai [7])
${\rm ex}(n,P_{k})\leq\frac{(k-2)n}{2}$, where $k\geq 2$.
Theorem 2 (Erdős-Gallai [7])
${\rm ex}(n,{\cal C}_{\geq k})\leq\frac{(k-1)(n-1)}{2}$, where $k\geq 3$.
If $k-1|n$, then the union of $n/(k-1)$ vertex disjoint $K_{k-1}$’s shows that the upper bound in Theorem 1 is tight. Similarly, if $k-2|n-1$, then the graph consists of $(n-1)/(k-2)$ copies of $K_{k-1}$ sharing a common vertex gives the tightness of the upper bound in Theorem 2. Additionally, one can see Theorem 1 is a direct consequence of Theorem 2.
For general $n$, Faudree and Schelp [12] and independently Kopylov [15] determined the value of ${\rm ex}(n,P_{k})$. Meanwhile, Woodall [24] obtained the value of ${\rm ex}(n,{\cal C}_{\geq k})$ and showed the result for ${\rm ex}(n,P_{k})$ is a corollary of the one for ${\rm ex}(n,{\cal C}_{\geq k})$.
In the study of Turán number of graphs, the celebrated Erdős-Stone-Simonovits Theorem [8, 9] gives an asymptotic formula for ${\rm ex}(n,{\cal H})$ when ${\cal H}$ contains no bipartite graphs:
${\rm ex}(n,{\cal H})=\left(1-\frac{1}{p({\cal H})}\right)\binom{n}{2}+o(n^{2}),$
where $p({\cal H})=\min\{\chi(H)-1:H\in{\cal H}\}$.
Many existing results study the Turán number of a single graph. For the case of ${\cal H}$ with two graphs, a classical result by Chvátal and Hanson [4] gives the Turán number of =${\cal H}=\{M_{\nu+1},K_{1,\Delta+1}\}$. A special case where $\nu=\Delta=k$ was first proved by Abbott, Hanson, and Sauer [1]. As a common generalization of Turán’s Theorem and Erdős-Gallai Theorem for the Turán number of matchings, Alon and Frankl [2] recently obtained the value of ${\rm ex}(n,{\cal H})$ for =${\cal H}=\{K_{r},M_{k}\}$. Wang, Hou, and Ma [23] proved a spectral analogue of this result. Taking a step forward, Katona and Xiao [14] investigated the Turán number of the family =${\cal H}=\{K_{r},P_{k}\}$. To state their results, we need to introduce the following operation of graphs. Given two vertex disjoint graphs $F_{1}$ and $F_{2}$, the join of $F_{1}\vee F_{2}$ is a graph obtained from $F_{1}\cup F_{2}$ by connecting each vertex of $F_{1}$ and each vertex of $F_{2}$. For $r\leq\lfloor k/2\rfloor$, Katona and Xiao proved the following result.
Theorem 3 (Katona-Xiao [14])
Let =${\cal H}=\{K_{r},P_{k}\}$. If $r\leq\lfloor k/2\rfloor$ and $n$ is large enough, then
${\rm ex}(n,{\cal H})={\rm ex}(\lfloor k/2\rfloor-1,K_{r-1})+(\lfloor k/2\rfloor-1)(n-\lfloor k/2\rfloor+1),$
and $T(\lfloor k/2\rfloor-1,r-2)\vee I_{n-\lfloor k/2\rfloor+1}$ is an extremal graph.
For $r\geq\lfloor k/2\rfloor+1$, they made the following conjecture.
Conjecture 1 (Katona-Xiao [14])
Assume $\lfloor k/2\rfloor+1\leq r<k$. If $k\geq 3$ is odd and $k-1|n$, then ${\rm ex}(n,{\cal H})=\tfrac{n}{k-1}{\rm ex}(k-1,K_{r})$ for sufficiently large $n$. If $k\geq 4$ is even, then ${\rm ex}(n,{\cal H})=\binom{\lfloor k/2\rfloor-1}{2}+(\lfloor k/2\rfloor-1)(n-\lfloor k/2\rfloor+1)$ for sufficiently large $n$.
In a recent paper, Fang, Zhu, and Chen [11] verified Conjecture 1. Katona and Xiao [14] also conjectured the asymptotic value of ${\rm ex}(n,\{P_{k},H\})$ for a nonbipartite graph $H$. This conjecture was proved by Liu and Kang [16].
As we mentioned above, the result for the Turán number of $P_{k}$ is an easy corollary of the one for the Turán number of ${\cal C}_{\geq k}$. In the same spirit, we study the Turán number of =${\cal H}=\{K_{r},{\cal C}_{\geq k}\}$ in this paper. If $r\geq k$, then we can see ${\rm ex}(n,{\cal H})={\rm ex}(n,{\cal C}_{\geq k})$. Moreover, if $n\leq k-1$, then ${\rm ex}(n,{\cal H})={\rm ex}(n,K_{r})$. Thus we assume
$3\leq r<k\leq n$ (1)
throughout this paper.
Assume that $n-1=p(k-2)+q$ with $q\leq k-3$. Define $F(n,k,r)$ as the graph which consists of $p$ copies of $T(k-1,r-1)$ and one copy of $T(q+1,r-1)$ sharing a common vertex. Let
$f(n,k,r)=p{\rm ex}(k-1,K_{r})+{\rm ex}(q+1,K_{r}).$
Note that $f(n,k,r)=e(F(n,k,r))$. Additionally, let
$G_{1}=K_{\lfloor(k-1)/2\rfloor}\vee I_{n-\lfloor(k-1)/2\rfloor}$
and
$G_{2}=T(\lfloor(k-1)/2\rfloor,r-2)\vee I_{n-\lfloor(k-1)/2\rfloor}.$
Our main results are the following ones.
Theorem 4
Let =${\cal H}=\{K_{r},{\cal C}_{\geq k}\}$ and $n-1=p(k-2)+q$. If $k\geq 6$ is even and $\lfloor(k-1)/2\rfloor+2\leq r<k\leq n$, then
${\rm ex}(n,{\cal H})=f(n,k,r).$
If $k\geq 5$ is odd and $\lfloor(k-1)/2\rfloor+2\leq r<k\leq n$, then
${\rm ex}(n,{\cal H})=\max\left\{f(n,k,r),e(G_{1})\right\}.$
Remark 1: For even $k$, if $k=4$, then $r=3$ by the lower bound assumption. Furthermore, the assumption $G$ being $\{K_{3},{\cal C}_{\geq 4}\}$-free implies that $G$ does not contain cycles. Thus $G$ either is a tree or is a forest. This case is trivial. As the assumption $3\leq r<k$, it is natural to assume $k\geq 5$ for odd $k$.
For $r\leq\lfloor(k-1)/2\rfloor+1$, we prove the following result.
Theorem 5
Let =${\cal H}=\{K_{r},{\cal C}_{\geq k}\}$. If $r\leq\lfloor(k-1)/2\rfloor+1$ and $n\geq k\geq 5$, then
${\rm ex}(n,{\cal H})=e(G_{2})$
provided $n\geq\tfrac{k^{3}}{4}$.
As an application of Theorem 4, we prove the following result.
Theorem 6
Let =${\cal H}=\{K_{r},P_{k}\}$ and $\lfloor k/2\rfloor+1\leq r<k\leq n$. Assume that $n=p(k-1)+q$. If $k\geq 3$ is odd, then
${\rm ex}(n,{\cal H})=p{\rm ex}(k-1,K_{r})+{\rm ex}(q,K_{r}).$
If $k\geq 4$ is even, then
${\rm ex}(n,{\cal H})=\max\left\{p{\rm ex}(k-1,K_{r})+{\rm ex}(q,K_{r}),\binom{\lfloor k/2\rfloor-1}{2}+(\lfloor k/2\rfloor-1)(n-\lfloor k/2\rfloor+1)\right\}.$
Apparently, for $k$ being odd and $k-1|n$, Theorem 6 gives that ${\rm ex}(n,{\cal H})=\tfrac{n}{k-1}{\rm ex}(k-1,K_{r})$ for all $n$. For $k$ being even and $k-1|n$, one can verify
$\frac{n}{k-1}{\rm ex}(k-1,K_{r})<\binom{\lfloor k/2\rfloor-1}{2}+(\lfloor k/2\rfloor-1)(n-\lfloor k/2\rfloor+1)$
for sufficiently large $n$ and then ${\rm ex}(n,{\cal H})=\binom{\lfloor k/2\rfloor-1}{2}+(\lfloor k/2\rfloor-1)(n-\lfloor k/2\rfloor+1)$ in this case. Therefore, we prove Conjecture 1 in a stronger form as Theorem 6 gives us the value of ${\rm ex}(n,{\cal H})$ for all $n$.
In order to prove Theorem 4 and Theorem 5, we recall the study of ${\rm ex}(n,{\cal C}_{\geq k})$. Actually, to determine the Turán number of ${\cal C}_{\geq k}$, essentially one need to consider the 2-connected analogue. To be precisely, let ${\rm ex}_{2\textrm{-conn}}(n,{\cal C}_{\geq k})$ be the maximum possible number of edges in a 2-connected $n$-vertex ${\cal C}_{\geq k}$-free graph.
Following the notation in the literature, for $2\leq a\leq\lfloor(k-1)/2\rfloor$, let $H(n,a,k)$ be a graph with vertex set $A\cup B\cup C$ such that $|A|=k-2a$, $|B|=a$, and $|C|=n-k+a$. Moreover, $A\cup B$ is a clique and each vertex in $B$ is adjacent to each vertex from $C$. Let $e(H(n,a,k))=h(n,a,k)=\binom{k-a}{2}+a(n-(k-a))$. In 1976, Woodall [24] conjectured that ${\rm ex}_{2\textrm{-conn}}(n,{\cal C}_{\geq k})=\max\{h(n,2,k),h(n,\lfloor(k-1)/2\rfloor,k)\}$ and verified this conjecture for large $n$. Fan, Lv, and Wang [10] proved the remaining case. Note that Kopylov resolved this conjecture completely in the paper [15].
For =${\cal H}=\{K_{r},{\cal C}_{\geq k}\}$, let ${\rm ex}_{2\textrm{-conn}}(n,{\cal H})$ be the maximum possible number of edges in a 2-connected $n$-vertex ${\cal H}$-free graph. For $2\leq a\leq\lfloor(k-1)/2\rfloor$, we vary the graph $H(n,a,k)$ through replacing the clique $K_{k-a}$ by the Turán graph $T(k-a,r-1)$ and the resulting graph is denoted by $G_{r}(n,a,k)$. Let
$g_{r}(n,a,k)=(n-k+a)a+{\rm ex}(k-a,K_{r}).$
Note that $g_{r}(n,a,k)=e(G_{r}(n,a,k))$ and ${\rm ex}(k-a,K_{r})=\binom{k-a}{2}$ for $r>k-a$. We will prove the following results for ${\rm ex}_{2\textrm{-conn}}(n,{\cal H})$.
Theorem 7
Let =${\cal H}=\{K_{r},{\cal C}_{\geq k}\}$. If $\lfloor(k-1)/2\rfloor+2\leq r<k$ and $n\geq k\geq 5$, then
${\rm ex}_{2\textrm{-conn}}(n,{\cal H})=\max\left\{g_{r}(n,2,k),g_{r}(n,\lfloor(k-1)/2\rfloor,k)\right\}.$
For $r\leq\lfloor(k-1)/2\rfloor+1$, our result is as follows.
Theorem 8
Let =${\cal H}=\{K_{r},{\cal C}_{\geq k}\}$. If $r\leq\lfloor(k-1)/2\rfloor+1$ and $n\geq k\geq 5$, then
${\rm ex}_{2\textrm{-conn}}(n,{\cal H})=e(G_{2})$
provided $n\geq\tfrac{k^{2}}{2}$.
Here if $k=4$, then $r=2$ and the problem is not interesting. Thus we assume $k\geq 5$.
Since proofs of results in this paper involve different techniques, we will present them one by one so that each of them will be self-contained. Namely, we organize the rest of this paper as follows. In Section 2, we will prove Theorem 7 and Theorem 8. Proofs of Theorem 4, Theorem 5, and Theorem 6 will be presented in Section 3. In Section 4, we will mention a few concluding remarks.
2 Proofs of Theorem 7 and Theorem 8
2.1 Proof of Theorem 7
The proof of Theorem 7 leverages on a variant of the method introduced by Kopylov [15]. Actually, the method introduced by Kopylov [15] turns out to be very useful in the study of Turán type problems for graphs with bounded circumference, for example [6, 17, 18, 20, 25].
We outline the proof of Theorem 7 as follows. Assume $G$ is an edge-maximal counterexample and $t=\lfloor(k-1)/2\rfloor$. Let $H(G,t)$ be the $(t+1)$-core of $G$. We can show that $H(G,t)$ is not empty. Furthermore, we prove that $H(G,t)$ either is a clique or is a $K_{r}$-saturated graph. For the former case, we utilize a variant of the method from [15] to show the existence of a cycle with length at least $k$, which is a contradiction to the assumption of $G$. For the latter case, the key new idea is that we can treat $H(G,t)$ as a clique (see Lemma 3) and show the contradiction by repeating the argument in the former case.
We first prove a number of lemmas and the proof of Theorem 7 will be presented at the end of Subsection 2.1. Readers can skip proofs of lemmas and go the proof of Theorem 7 directly.
For the rest of this subsection, we fix
$t=\lfloor(k-1)/2\rfloor.$
Recall the definition $g_{r}(n,k,a)=(n-k+a)a+{\rm ex}(k-a,K_{r}).$ We next show that $g_{r}$ is a convex function with $a$ being the independent variable.
Lemma 1
The function $g_{r}(n,a,k)$ is convex with $a$ being the independent variable, provided $r\geq 3$ and $2\leq a\leq t$.
Proof: Let $m(a)=g_{r}(n,a+1,k)-g_{r}(n,a,k)$. Then
$m(a)$ $=g_{r}(n,a+1,k)-g_{r}(n,a,k)$
$=(n-k+a+1)(a+1)+{\rm ex}(k-a-1,K_{r})-(n-k+a)a-{\rm ex}(k-a,K_{r})$
$=n-k+2a+1+\big{(}{\rm ex}(k-a-1,K_{r})-{\rm ex}(k-a,K_{r})\big{)}.$
To get $T(k-a,r-1)$, one can add a new vertex to a smallest partite of $T(k-a-1,r-1)$ and connect this vertex to all vertices from other parties. Thus
${\rm ex}(k-a,K_{r})-{\rm ex}(k-a-1,K_{r})=k-a-1-\lfloor(k-a-1)/(r-1)\rfloor.$
It follows that
$m(a+1)-m(a)$ $=n-k+2(a+1)+1+\big{(}{\rm ex}(k-a-2,K_{r})-{\rm ex}(k-a-1,K_{r})\big{)}$
$\quad-(n-k+2a+1)-\big{(}{\rm ex}(k-a-1,K_{r})-{\rm ex}(k-a,K_{r})\big{)}$
$=2-(k-a-2-\lfloor(k-a-2)/(r-1)\rfloor)+(k-a-1-\lfloor(k-a-1)(r-1)\rfloor)$
$=3-\lfloor(k-a-1)/(r-1)\rfloor+\lfloor(k-a-2)/(r-1)\rfloor$
$\geq 3-(k-a-1)/(r-1)+(k-a-2)/(r-1)-1$
$=2-\frac{1}{r-1}>0.$
Thus $g_{r}(n,k,a)$ is convex if $r\geq 3$ and $2\leq a\leq t$. $\square$
We will need the following two results on the length of a longest cycle in a graph.
Theorem 9 (Dirac [5])
Let $G$ be a 2-connected graph of order $n$ with minimum degree $\delta$. Then $c(G)\geq\mbox{min}\{2\delta,n\}$, where $c(G)$ is the length of a longest cycle in $G$.
Lemma 2 (Kopylov [15])
Let $G$ be a $2$-connected $n$-vertex graph with a path $P$ of $m$ edges with endpoints $x$ and $y$. For $v\in V(G)$, let $d_{P}(v)=|N(v)\cap V(P)|$. Then $G$ contains a cycle of length at least $\min\{m+1,d_{P}(x)+d_{P}(y)\}.$
A graph $G$ is $K_{r}$-saturated if $G$ is $K_{r}$-free and adding any nonedge to $G$ will create a copy of $K_{r}$. In other words, any two nonadjacent vertices are in a $K_{r}-e$, here $K_{r}-e$ is the clique $K_{r}$ with one edge removed. The next lemma plays an important role in our proof of Theorem 7.
Lemma 3
Assume that $H$ is an induced subgraph of $G$ and $H$ is $K_{r}$-saturated with $\delta(H)\geq\lfloor(k-1)/2\rfloor+1$ and $k\geq 5$. Let $G^{\prime}$ be a graph by adding all nonedges with endpoints in $H$. If $r\geq\lfloor(k-1)/2\rfloor+2$ and the length of a longest cycle in $G^{\prime}$ is at least $k$, then there is a cycle of length at least $k$ in $G$.
Proof: Let $\overline{H}$ be the complement graph of $H$. Among all cycles of length at least $k$ in $G^{\prime}$, we pick $C$ as the one satisfies the following two conditions:
(a): $|E(C)\cap E(\overline{H})|$ is minimized; and
(b): the length is maximized.
Let $C=a_{1}a_{2}\cdots a_{\ell}a_{1}$ with $\ell\geq k$ and vertices of $C$ are ordered clockwisely. For each $1\leq i\leq\ell$, we will write $a_{i}^{+}$ for $a_{i+1}$ and $a_{i}^{-}$ for $a_{i-1}$ in some circumstances, here the addition is under modulo $\ell$. We claim that $C$ is a cycle in $G$. If $C$ contains edges from $E(\overline{H})$, then we next find a contradiction to the definition of $C$. By the construction of $G^{\prime}$, we may assume that $a_{1}a_{\ell}\in E(\overline{H})$. Note that $\{a_{1},a_{\ell}\}\subset V(H)$. Define $N_{H}(a)=N_{G}(a)\cap V(H)$. Assume that $N_{C}(a_{1})\cap N_{C}(a_{\ell})=\{a_{i_{1}},\ldots,a_{i_{q}}\}$ with vertices ordered clockwisely. Let $P_{1}=a_{i_{1}}a_{i_{1}}^{+}\cdots a_{i_{q}}$, $P_{2}=a_{1}a_{2}\cdots a_{i_{q}}$, and $P_{3}=a_{i_{1}}a_{i_{1}}^{+}\cdots a_{\ell}$.
Claim 1
$(1)$ $N_{G}(a_{1})\cap N_{G}(a_{\ell})\subset V(C)$, $N_{H}(a_{1})\subset V(C)$, and $N_{H}(a_{\ell})\subset V(C)$.
$(2)$ For each $11\leq j\leq q-$ 1, $a_{i_{j}}^{+}\neq a_{i_{j+1}}$.
$(3)$ $a_{1}a_{i_{j}}^{+}\not\in E(G)$ for each $1\leq j\leq q$ and $a_{\ell}a_{i_{j}}^{-}\not\in E(G)$ for each $1\leq j\leq q$.
$(4)$ $|V(P_{1})|\in\{k-3,k-2\},|V(P_{2})|\leq k-1$, and $|V(P_{3})|\leq k-1.$
Proof of Claim 1: For part $(1)$, if there is a vertex $v\in N_{G}(a_{1})\cap N_{G}(a_{\ell})$ such that $v\not\in V(C)$, then we get a new cycle $C_{1}=va_{1}\cdots a_{\ell-1}a_{\ell}v$ of length at least $k$ and $|E(C_{1})\cap E(\overline{H})|<|E(C)\cap E(\overline{H})|$. This is a contradiction to the choice of $C$. Therefore, $N_{G}(a_{1})\cap N_{G}(a_{\ell})\subset V(C)$. If there is a vertex $v\in N_{H}(a_{1})$ such that $v\not\in V(C)$, As $H$ is a clique in $G^{\prime}$ and $v,a_{\ell}\in V(H)$, then $va_{1}\cdots a_{\ell}v$ is a cycle with at least $k$ edges in $G^{\prime}$. If $va_{\ell}\in E(G)$, then it violates the condition (a). If $va_{\ell}\not\in E(G)$, then it violates the condition (b). Thus $N_{H}(a_{1})\subset V(C)$. By symmetry, we also have $N_{H}(a_{\ell})\subset V(C)$.
For part $(2)$, if there is a $a_{i_{j}}$ such that $a_{i_{j}}^{+}=a_{i_{j+1}}$, then we get a new cycle $C_{1}=a_{1}a_{2}\cdots a_{i_{j}}a_{\ell}a_{\ell}^{-}\cdots a_{i_{j+1}}a_{1}$ with at least $k$ edges and $|E(C_{1})\cap E(\overline{H})|<|E(C)\cap E(\overline{H})|$. This is a contradiction to the definition of $C$.
The proof of part $(3)$ is the same as the one for part $(2)$ and it is skipped here.
For part $(4)$, we first claim $|V(P_{1})|\in\{k-3,k-2\}$. If $|V(P_{1})|\geq k-1$, then $C_{1}=a_{1}a_{2}\ldots a_{i_{q}}a_{1}$ is a new cycle with at least $k$ edges and $|E(C_{1})\cap E(\overline{H})|<|E(C)\cap E(\overline{H})|$, here note that $a_{1}a_{i_{q}}\in E(G)$. This is a contradiction to the definition of $C$. Next we claim $|V(P_{1})|\geq k-3.$ Recall $N_{C}(a_{1})\cap N_{C}(a_{\ell})=\{a_{i_{1}},\ldots,a_{i_{q}}\}$. By part $(1)$, we have $N_{G}(a_{1})\cap N_{G}(a_{\ell})\subset V(C)$. As $H$ is $K_{r}$-saturated, $\{a_{1},a_{\ell}\}\subset V(H)$, and $a_{1}$ is not adjacent to $a_{\ell}$, we get
$q=|N_{G}(a_{1})\cap N_{G}(a_{\ell})|\geq r-2.$
By part $(2)$, we have $a_{i_{j}}^{+}\neq a_{i_{j+1}}$, i.e., $i_{j+1}-i_{j}\geq 2$. One can observe that
$V(P_{1})\geq 2(q-1)+1.$
Recall that $q\geq r-2\geq\lfloor(k-1)/2\rfloor$. It follows that $V(P_{1})\geq 2\lfloor(k-1)/2\rfloor-1\geq k-3.$ Thus $|V(P_{1})|\in\{k-3,k-2\}.$ Finally, we claim $|V(P_{2})|\leq k-1$ and $|V(P_{3})|\leq k-1$. By symmetry, one can suppose that $|V(P_{2})|\geq k$. Now $C_{1}=a_{1}a_{2}\cdots a_{i_{q}}a_{1}$ is a new cycle with at least $k$ edges and $|E(C_{1})\cap E(\overline{H})|<|E(C)\cap E(\overline{H})|$. This is a contradiction to the definition of $C$. $\square$
We continue to prove the lemma. For the case of $r\geq\lfloor(k-1)/2\rfloor+3$, it follows that $q\geq r-2\geq\lfloor(k-1)/2\rfloor+1$ and $V(P_{1})\geq 2\lfloor(k-1)/2\rfloor+1\geq k-1.$ This is a contradiction to part $(4)$ of Claim 1. It remains to show the desired contradiction for the case where $r=\lfloor(k-1)/2\rfloor+2$. Repeating the argument above, we can observe that $q=\lfloor(k-1)/2\rfloor$. Notice that $|V(P_{1})|\in\{k-3,k-2\}$ by part $(4)$ of Claim 1.
If $|V(P_{1})|=k-2$, then $\ell=k$, $a_{i_{1}}=a_{2}$, and $a_{i_{q}}=a_{k-1}$ by $(4)$ of Claim 1. Recall $|V(P_{1})|\geq 2(q-1)+1$. For the case that $k$ is odd, then $|V(P_{1})|\geq k-2$. This implies $a_{i_{j+1}}=a_{i_{j}}+2$ for any $11\leq i\leq q-$ 1 . Since $|N_{H}(a_{1})|\geq\lfloor(k-1)/2\rfloor+1$, $N_{H}(a_{1})\subset V(C)$ by part $(1)$ of Claim 1, and $|N_{G}(a_{1})\cap N_{G}(a_{\ell})|=q=\lfloor(k-1)/2\rfloor$, it follows that there is some $a_{i}$ such that $a_{i}\in N_{H}(a_{1})$ but $a_{i}\not\in N_{G}(a_{1})\cap N_{G}(a_{\ell})$. The only possibility is that $a_{i}=a_{i_{j}}^{+}$ for some $a_{i_{j}}$. However, this is a contradiction to part $(3)$ of Claim 1. For the case that $k$ is even, there exists a unique $1\leq j\leq q$ such that $a_{i_{j+1}}=a_{i_{j}}+3$. Similarly, there are $a_{i}$ and $a_{j}$ such that $a_{i}\in N_{H}(a_{1})\setminus N_{G}(a_{\ell})$ and $a_{j}\in N_{H}(a_{\ell})\setminus N_{G}(a_{1})$. Part $(3)$ of Claim 1 yields that the only possibility is $a_{i}=a_{i_{j+1}}^{-}$ and $a_{j}=a_{i_{j}}^{+}$. Now we can get a new cycle $C_{1}=a_{1}a_{i_{j+1}}^{-}a_{i_{j+1}}\cdots a_{\ell}a_{i_{j}}^{+}a_{i_{j}}\cdots a_{1}$ with at least $k$ edges and $|E(C_{1})\cap E(\overline{H})|<|E(C)\cap E(\overline{H})|$. This is a contradiction to the definition of $C$.
For $|V(P_{1})|=k-3$, we get that $k$ is even and $a_{i_{j+1}}=a_{i_{j}}+2$ for any $11\leq i\leq q-$ 1. Recall part $(4)$ of Claim 1 and the assumption $|V(C)|\geq k$. Either $\ell=k,a_{i_{1}}=a_{3},a_{i_{q}}=k-1$, or $\ell=k,a_{i_{1}}=a_{2},a_{i_{q}}=k-2$, or $\ell=k+1,a_{i_{1}}=a_{3},a_{i_{q}}=k-1.$ For the first case, there is a vertex $a_{i}\in N_{H}(a_{\ell})\setminus N_{G}(a_{1})$ as the minimum degree assumption. It is only possible that $a_{i}=a_{i_{j}}^{-}$ for some $a_{i_{j}}$. This is a contradiction to part $(2)$ of Claim 1. One can show the same contradiction for the second case. For the third case, consider the cycle $C_{1}=a_{1}\cdots a_{i_{1}}a_{\ell}\cdots a_{i_{2}}a_{1}$. Note that $C_{1}$ contains $k$ edges and $|E(C_{1})\cap E(\overline{H})|<|E(C)\cap E(\overline{H})|$. This is a contradiction to the definition of $C$. The lemma is proved. $\square$
We are ready to prove Theorem 7.
Proof of Theorem 7: Notice that $r\geq t+2$. For the lower bound, we recall the definition of $G_{r}(n,a,k)$. It is obvious that $G_{r}(n,a,k)$ is $\{K_{r},{\cal C}_{\geq k}\}$-free and the lower bound follows. We mention a special case where $k$ is even, $r=t+2$, and $a=t=k/2-1$. In this case, the subgraph of $G_{r}(n,a,k)$ induced by $A\cup B$ is $K_{r}$ with one edge removed.
For the upper bound, if the assertion in Theorem 7 is not true, then let $G$ be an edge-maximal counterexample. That is $G$ is ${\cal H}$-free and
$e(G)>\max\{g_{r}(n,2,k),g_{r}(n,t,k)\}.$
Moreover, adding any nonedge to $G$ either creates a $K_{r}$ or results a cycle with length at least $k$.
Let $H(G,t)$ be the $(t+1)$-core of $G$, i.e., the largest induced subgraph of $G$ with minimum degree at least $t+1$.
Claim 2
$H(G,t)$ is not empty.
Proof: Note that if $k$ is odd, then the subgraph of $G_{r}(n,t,k)$ induced by $A\cup B$ is a clique of size $t+1$ and each vertex in $C$ is adjacent to all vertices in $B$. If $k$ is even, then the subgraph induced by $A\cup B$ contains $K_{t+1}$ with one edge removed as a subgraph. Therefore, $e(G_{r}(n,t,k))=g_{r}(n,t,k)\geq(n-t)t+\binom{t}{2}$. If $H(G,t)$ is empty, then in the process of defining $H(G,t)$, we remove at most $t$ edges for each of the first $n-t$ vertices and the subgraph of the last $t$ vertices contains at most $\binom{t}{2}$ edges. Thus $e(G)\leq t(n-t)+\binom{t}{2}$ and this is a contradiction to the lower bound for $e(G)$. $\square$
Claim 3
If $x,y\in H(G,t)$ and $x$ is not adjacent to $y$, then there is no $xy$-path with at least $k-1$ edges in $G$.
Proof: If the assertion does not hold, then we choose vertices $x$ and $y$ from $H(G,t)$ such that a longest path in $G$ from $x$ to $y$ contains the largest number of edges among all such nonadjacent pairs. Let $P$ be a longest $xy$-path in $G$. We assert that all neighbors of $x$ in $H(G,t)$ lie in $P$. If $x$ has a neighbor $x^{\prime}\in H(G,t)\setminus P$, then either $x^{\prime}y\in E(G)$ or $x^{\prime}y\notin E(G)$. In the former case, $x^{\prime}xPyx^{\prime}$ is a cycle of length at least $k$ in $G$, a contradiction to the assumption for $G$. For the latter case, $x^{\prime}xPy$ is a longer path, which is a contradiction to the selection of $x$ and $y$. There is a contradiction in each case. We can show the same assertion for $y$ similarly. Therefore, $d_{P}(x)\geq t+1$ and $d_{P}(y)\geq t+1$. By Lemma 2, $G$ contains a cycle of length at least $\min\{k,d_{P}(x)+d_{P}(y)\}=k$, a contradiction. $\square$
The next claim follows from the assumption for $G$.
Claim 4
$H(G,t)$ is either a clique or a $K_{r}$-saturated graph.
Claim 5
$t+2\leq h=|H(G,t)|\leq k-2.$
Proof: The lower bound is trivial as the minimum degree assumption. Note that $r\geq 4$ because we assume $k\geq 5$ and $r\geq t+2$. As $H(G,t)$ is either a clique or $K_{r}$-saturated, then it is 2-connected. If $h\geq k$, then Theorem 9 ensures that $H(G,t)$ contains a cycle of length at least $\min\{h,2(t+1)\}\geq k$, a contradiction. If $h=k-1$, then a classical result by Ore [21] implies that $H(G,t)$ is Hamiltonian connected . Since $H(G,t)$ is an induced subgraph of a 2-connected graph $G$, for any vertex $x$ not in $H(G,t)$, there are two internally disjoint paths from $x$ to $H(G,t)$. Let $y$ and $z$ be the endpoints of these two paths in $H(G,t)$. As there is a Hamilton path from $y$ to $z$ in $H(G,t)$, we get that $G$ contains a cycle with at least $k$ edges, a contradiction. $\square$
Let $H(G,k-h)$ be the $(k-h+1)$-core of $G$.
Claim 6
$H(G,t)\neq H(G,k-h)$.
Proof: Observe that $2\leq k-h\leq t$. If $H(G,t)=H(G,k-h)$, then in the process of defining $H(G,t)$, we remove at most $k-h$ edges for the first $n-h$ vertices and the subgraph in the last $h$ vertices contains at most ${\rm ex}(h,K_{r})$ edges as $G$ is $K_{r}$-free. We get that
$e(G)\leq{\rm ex}(h,K_{r})+(n-h)(k-h)=g_{r}(n,k-h,k)\leq\max\{g_{r}(n,2,k),g_{r}(n,t,k)\},$
a contradiction. Here the last inequality holds as the range of $k-h$ and Lemma 1. $\square$
We first consider the case where $H(G,t)$ is not a clique, i.e., $H(G,t)$ is $K_{r}$-saturated. Denote $X=V(H(G,t))$ and $Y=V(H(G,k-h))$. As we already proved $H(G,t)\neq H(G,k-h)$, it follows that $X$ is a proper subset of $Y$. In addition, for each $y\in Y\setminus X$, there is $x\in X$ such that $x$ and $y$ are not adjacent in $G$. Otherwise, $d_{G}(y)\geq h\geq t+2$ and $y$ should be a vertex in $X$, a contradiction.
If $r\geq t+3$, then for any nonadjacent $x$ and $y$ with $x\in X$ and $y\in Y\setminus X$, we claim that adding the edge $xy$ to $G$ must create a cycle with length at least $k$. Otherwise, $x$ and $y$ are contained in $K_{r}-e$ by the assumption of $G$. As $r\geq t+3$, all vertices in $K_{r}-e$ has at least $t+1$ neighbors in $K_{r}-e$. Then $y\in H(G,t)$, a contradiction. Let
$Q=\{(x,y):x\in X,y\in Y\setminus X,\textrm{ and }x\textrm{ is not adjacent to }y\}.$
Let $G^{\prime}$ be a new graph by adding all nonedges in $X$. For each pair $(x,y)\in Q$, we already show that there is an $xy$-path with at least $k-1$ edges in $G$. Then the same observation also holds in $G^{\prime}$. We choose the pair $(x,y)\in Q$ such that a longest path $xy$-path, say $P_{xy}$, in $G^{\prime}$ contains the largest number of edges among all pairs in $Q$. Notice that $N_{G^{\prime}}(y)=N_{G}(y)$. We claim that $N_{G^{\prime}}(y)\cap Y\subset V(P_{xy})$. If not, then let $z\in N_{G^{\prime}}(y)\cap Y$ and $z\not\in V(P_{xy})$. As the choice of the pair $(x,y)$, we get that $x$ and $z$ must be adjacent. Thus $xP_{xy}yzx$ is a cycle of length at least $k$ in $G^{\prime}$. By Lemma 3, $G$ contains a cycle of length $k$, a contradiction. Similarly, we can show $X\subset V(P_{xy})$. Therefore, by Lemma 2, $G^{\prime}$ contains a cycle of length at least $\min\{k,h-1+k-h+1\}=k$, here note that $X$ is a clique in $G^{\prime}$. Then $G$ contains a cycle with at least $k$ edges by Lemma 3. We obtained a contradiction in each case.
For $r=t+2$, if there are vertices $x\in X$ and $y\in Y\setminus X$ such that $xy\not\in E(G)$ and adding the edge $xy$ to $G$ creates a cycle with length at least $k$, then one can obtain the same contradiction by reusing the argument above. Therefore, for any nonadjacent vertices $x$ and $y$ such that $x\in X$ and $y\in Y\setminus X$, adding the edge $xy$ to $G$ must create a copy of $K_{t+2}$. It yields that $d_{Y}(y)\geq t$. We next show $Y\setminus X$ is an independent set. It only needs to consider the case of $|Y\setminus X|\geq 2$. Suppose that there is an edge $y_{1}y_{2}$ with $\{y_{1},y_{2}\}\subset Y\setminus X$. For $i=1,2$, let $x_{i}\in X$ be a vertex such that $x_{i}$ and $y_{i}$ are not adjacent. Additionally, by the assumption, there is a $t$-set $T_{i}\subset Y$ such that $\{x_{i},y_{i}\}\cup T_{i}$ induces a $K_{t+2}-e$. Observe that the subgraph induced by $\cup_{i=1}^{2}\{x_{i},y_{i}\}\cup T_{1}\cup T_{2}$ has minimum degree $t+1$. Thus $y_{i}\in X$. This is a contradiction to the choice of $y_{i}$ and then $Y\setminus X$ is an independent set. Fix nonadjacent vertices $y\in Y\setminus X$ and $x\in X$. Let $T\subset Y$ be the vertex of a $K_{t}$ such that $\{x,y\}\cup T$ induces $K_{t+2}-e$. As $Y\setminus X$ is an independent set, we get that $T\subset X$. Moreover, $N_{Y}(y)=T$ and $y$ is not adjacent to each $x^{\prime}\in X\setminus T$. Otherwise, $y$ has at least $t+1$ neighbors in $X$ and $y$ should be a vertex in $X$, a contradiction. Therefore, $\{x^{\prime},y\}\cup T$ is $K_{t+2}-e$ for any $x^{\prime}\in X\setminus T$. Since $G$ is $K_{t+2}$-free, $X\setminus T$ is an independent set. This implies that $d_{X}(x^{\prime})=t$ for each $x^{\prime}\in X\setminus T$, a contradiction to the definition of $H(G,t)$.
Next, we consider the case where $H(G,t)$ is a clique. If $r\geq t+3$, then we repeat the argument above by defining $G^{\prime}=G$ and get a contradiction. For $r=t+2$, since $t+2\leq|H(G,t)|\leq k-2$, then $H(G,t)$ contains a $K_{r}$, a contradiction. The proof of Theorem 7 is complete. $\square$
2.2 Proof of Theorem 8
As we are not able to prove Lemma 3 under assumptions in Theorem 8, we apply a different approach to prove Theorem 8. Namely, let $G$ be a 2-connected graph which is both $K_{r}$-free and ${\cal C}_{\geq k}$-free. A longest cycle in $G$ is denoted by $C$. Let $e(C)$ be the number of edges contained in $C$ and $e(G-C)+e(G-C,C)$ be the number of edges which has at most one endpoint in $C$. Apparently,
$e(G)=e(C)+e(G-C)+e(G-C,C).$
We will upper bound $e(C)$ and $e(G-C)+e(G-C,C)$ separately. Lemma 9 will give us an upper bound for $e(G)$.
We start with the following estimate for the number of edges in $C$.
Lemma 4
Let $G$ be a $K_{r}$-free graph and $C$ be a longest cycle in $G$ of length $c$, where $r\leq\lfloor c/2\rfloor+1$. If there exists a vertex $u\in V(G-C)$ with $d_{C}(u)=\lfloor c/2\rfloor$, then
$e(C)\leq{\rm ex}(\lfloor c/2\rfloor,K_{r-1})+\lfloor c/2\rfloor\lceil c/2\rceil.$
Proof: If $c$ is even, then let $C=a_{1}b_{1}a_{2}b_{2}\cdots a_{c/2}b_{c/2}a_{1}$ be the longest cycle. Since $d_{C}(u)=c/2$, without loss of generality, assume $A=V(C)\cap N_{C}(u)=\{a_{1},\ldots,a_{i},\ldots,a_{c/2}\}$. Let $B=\{b_{1},\ldots,b_{i},\ldots,b_{c/2}\}$. Then $e(B)=0$. Otherwise, there is a longer cycle by including $u$, a contradiction. As $G[C]$ is also $K_{r}$-free, it follows that
$e(C)$ $=e(A)+e(B)+e(A,B)$
$\leq{\rm ex}(c/2,K_{r-1})+(c/2)^{2}$
$={\rm ex}(c/2,K_{r-1})+(c/2)^{2}.$
If $c$ is odd, then let $C=a_{1}b_{1}a_{2}b_{2}\cdots a_{\lfloor c/2\rfloor}b_{\lfloor c/2\rfloor}b_{\lceil c/2\rceil}a_{1}$ be the longest cycle. Similarly, we assume $A=V(C)\cap N_{C}(u)=\{a_{1},\ldots,a_{i},\ldots,a_{\lfloor c/2\rfloor}\}$ and $B=\{b_{1},\ldots,b_{i},\ldots,b_{\lfloor c/2\rfloor},b_{\lceil c/2\rceil}\}$. Then $G[B]$ contains at most one edge, that is $b_{\lfloor c/2\rfloor}b_{\lceil c/2\rceil}$. Otherwise, there is a longer cycle by including $u$, a contradiction. Thus
$e(C)$ $=e(A)+e(B)+e(A,B)$
$\leq{\rm ex}(\lfloor c/2\rfloor,K_{r-1})+1+\lfloor c/2\rfloor\lceil c/2\rceil.$
The inequality above must be strict as $r\leq\lfloor c/2\rfloor+1$. Otherwise, a $K_{r-2}$ in $A$ together with vertices $b_{\lfloor c/2\rfloor}$ and $b_{\lceil c/2\rceil}$ form a $K_{r}$, a contradiction. Thus
$e(C)\leq{\rm ex}(\lfloor c/2\rfloor,K_{r-1})+\lfloor c/2\rfloor\lceil c/2\rceil$
in this case. The lemma is proved. $\square$
Let us recall the following classical result by Bondy.
Lemma 5 (Bondy [3])
Let $G$ be a graph on $n$ vertices and $C$ be a longest cycle of $G$ with order $c$. Then
$e(G-C)+e(G-C,C)\leq\lfloor c/2\rfloor(n-c).$
Woodall [24] made a conjecture on the maximum number of edges in a 2-connected graph which is ${\cal C}_{\geq k}$-free and has minimum degree at least $k\geq 2$. Ma and Ning [19] proved a stability result of Woodall’s conjecture. Previously, Füredi, Kostochka, and Verstraëte [13] obtained a stability result for Erdős-Gallai theorems on paths and cycles (see Theorem 1 and Theorem 2). In [19], Ma and Ning discovered the stability of Woodall’s conjecture is directly related to the above classical result of Bondy. Indeed, they proved a stability result of Lemma 5, see Theorem 3.2 in [19]. For our problem, we need the following variant of Theorem 3.2 in [19].
Lemma 6
Let $G$ be a graph on $n$ vertices and $C$ be a longest cycle in $G$ of length $c$, where $14\leq c\leq n-$ 1. If $e(G-C)+e(G-C,C)>(\lfloor c/2\rfloor-\tfrac{1}{2})(n-c)$, then one of the following holds:
$(1)$ there exists a vertex $u\in V(G-C)$ with $d_{C}(u)=\lfloor c/2\rfloor;$
$(2)$ there exists a cycle $C^{\prime}$in $G$ such that $|V(C\cap C^{\prime})\leq 1$. Moreover, if $V(C\cap C^{\prime})=\emptyset$, then $|C^{\prime}|\geq 2\lfloor\frac{c}{2}\rfloor-3.$ If $|V(C\cap C^{\prime})|=1$, then $|C^{\prime}|\geq 2\lfloor\frac{c}{2}\rfloor-1.$
The proof of the lemma above is similar to the one for Theorem $3.2$ in Ma and Ning’s paper [19] and it is omitted here. If $G$ is 2-connected, then Ma and Ning pointed out that $(2)$ does not occur. Thus the combination of Lemma 5 and Lemma 6 gives the following lemma.
Lemma 7
Let $G$ be a $2$-connected graph on $n$ vertices and $C$ be a longest cycle in $G$ of length $c$ with $14\leq c\leq n-$ 1.
$(1)$ If there exists a vertex $u\in V(G-C)$ with $d_{C}(u)=\lfloor c/2\rfloor$, then
$e(G-C)+e(G-C,C)\leq\lfloor c/2\rfloor(n-c).$
$(2)$ If $d_{C}(u)<\lfloor c/2\rfloor$ for each vertex $u\in V(G-C)$, then
$e(G-C)+e(G-C,C)\leq(\lfloor c/2\rfloor-\tfrac{1}{2})(n-c).$
Let $t=\lfloor(k-1)/2\rfloor$. The following inequality will be used in the proof of Lemma 9.
Lemma 8
If $k\leq n$ and $r\geq 3$, then
${\rm ex}(k-2,K_{r})+\left(t-\frac{1}{2}\right)(n-k+2)<{\rm ex}(k-1,K_{r})+\left(t-\frac{1}{2}\right)(n-k+1).$
Proof: Notice that ${\rm ex}(k-1,K_{r})$ and ${\rm ex}(k-2,K_{r})$ are the number of edges in the Turán graph $T(k-1,r-1)$ and $T(k-2,K_{r})$ respectively. The Turán graph $T(k-2,r-1)$ can be obtained by deleting a vertex from the largest partite of graph $T(k-1,r-1)$. This implies ${\rm ex}(k-1,K_{r})-{\rm ex}(k-2,K_{r})=k-1-\lceil(k-1)/(r-1)\rceil$. Thus it is sufficient to show
$k-1-\lceil(k-1)/(r-1)\rceil>t-1/2=\lfloor(k-1)/2\rfloor-1/2.$
As $r\geq 3$, then $k-1-\lceil(k-1)/(r-1)\rceil\geq k-1-\lceil(k-1)/2\rceil=\lfloor(k-1)/2\rfloor$. The inequality above is true. $\square$
The next lemma gives us an upper bound for the number of edges in a 2-connected $\{K_{r},{\cal C}_{\geq k}\}$-free graph which is of independent interest.
Lemma 9
Assume that $n\geq k\geq 5$ and $r\leq t+1$. Let $G$ be a $2$-connected $\{K_{r},{\cal C}_{\geq k}\}$-free graph with $n$ vertices and $C$ be a longest cycle in $G$.
$(1)$ If $|V(C)|=k-1$ and there exists $u\in V(G)\setminus V(C)$ such that $u$ has $t$ neighbors in $C$, then
$e(G)\leq{\rm ex}(t,K_{r-1})+t(n-t).$
$(2)$ If $k$ is even, $|V(C)|=k-2$, and there exists $u\in V(G)\setminus V(C)$ such that $u$ has $t$ neighbors in $C$, then
$e(G)\leq{\rm ex}(t,K_{r-1})+t(n-t).$
$(3)$ Otherwise, we have
$e(G)\leq\max\{(t-1/2)(n-1),{\rm ex}(k-1,K_{r})+(t-1/2)(n-k+1)\}.$
Proof: Note that $G$ is a 2-connected $n$-vertex graph that is $\{K_{r},{\cal C}_{\geq k}\}$-free. In addition, $C$ is a longest cycle in $G$. It implies that $14\leq c\leq k-1\leq n-$ 1. The lower bound follows from the assumption that $G$ is 2-connected.
We start to prove part $(1)$. Notice that $|V(C)|=k-1$ and there is a vertex $u\in V(G\setminus C)$ such that $d_{C}(u)=t$. By Lemma 7, we have
$e(G-C)+e(G-C,C)\leq t(n-k+1).$
Lemma 4 tells us that
$e(C)+e(G-C)+e(G-C,C)$ $\leq{\rm ex}(t,K_{r-1})+t(k-1-t)+t(n-k+1)$
$={\rm ex}(t,K_{r-1})+t(n-t).$
Thus the assertion in $(1)$ follows.
For part $(2)$, notice that $t=(k-2)/2$ as $k$ is even. Repeating the proof for part $(1)$, we can show part $(2)$.
For part $(3)$, if $|V(C)|=k-1$, then each $u\in V(G\setminus C)$ has at most $t-1$ neighbors in $C$. By Lemma 7, we have
$e(G-C)+e(G-C,C)\leq\left(t-\tfrac{1}{2}\right)(n-k+1).$
Apparently, $e(C)\leq{\rm ex}(k-1,K_{r})$. Thus
$e(G[C])+e(G-C)+e(G-C,C)\leq{\rm ex}(k-1,K_{r})+\left(t-\tfrac{1}{2}\right)(n-k+1).\$
If $|V(C)|\leq k-3$, then Theorem 2 yields that
$e(G)\leq\frac{k-3}{2}(n-1)\leq(t-\frac{1}{2})(n-1)$
no matter the parity of $k$. If $|V(C)|=k-2$ and $k$ is odd, then $e(G)\leq(t-\tfrac{1}{2})(n-1)$ also follows from Theorem 2. If $|V(C)|=k-2$ and $k$ is even, then each vertex from $u\in V(G\setminus C)$ has at most $t-1$ neighbors in $C$. By Lemma 7, we get that
$e(G-C)+e(G-C,C)\leq(t-\tfrac{1}{2})(n-k+2).$
Then $e(G)$ can be upper bounded as follows:
$e(G)$ $=e(G[C])+e(G-C)+e(G-C,C)$
$\leq{\rm ex}(k-2,K_{r})+(t-\tfrac{1}{2})(n-k+2)$
$<{\rm ex}(k-1,K_{r})+(t-\frac{1}{2})(n-k+1),$
the last inequality is true by Lemma 8. Part $(3)$ is proved. $\square$
We are now ready to prove Theorem 8.
Proof of Theorem 8: For the lower bound, recall that $t=\lfloor(k-1)/2\rfloor$ and the graph $G_{2}$. Observe that the clique number of $G_{2}$ is $r-1$ and a longest cycle has $12t\leq k-$ 1 edges. Moreover, this graph is 2-connected. The lower bound is proved.
For the upper bound, let $G$ be an $n$-vertex 2-connected graph which is $\{K_{r},{\cal C}_{\geq k}\}$-free. By Lemma 9, either
$e(G)\leq{\rm ex}(t,K_{r-1})+t(n-t).$
or
$e(G)\leq\max\{(t-1/2)(n-1),{\rm ex}(k-1,K_{r})+(t-1/2)(n-k+1)\}.$
We need only to consider the latter case. Observe that
$e(G)$ $\leq\max\{(t-1/2)(n-1),{\rm ex}(k-1,K_{r})+(t-1/2)(n-k+1)\}$
$\leq{\rm ex}(t,K_{r-1})+t(n-t)$
as $n\geq\tfrac{k^{2}}{2}$. The proof for Theorem 8 is complete. $\square$
3 Proofs of Theorem 4, Theorem 5, and Theorem 6
3.1 Proof of Theorem 6
We first present the proof of Theorem 6 as it follows from a well-known trick.
Proof of Theorem 6: Note that $n=p(k-1)+q$. For the lower bound, let $G_{3}$ be the graph which consists of $p$ copies of $T(k-1,r-1)$ and one copy of $T(q,r-1)$. In addition, let $G_{4}=K_{\lfloor k/2\rfloor-1}\vee I_{n-\lfloor k/2\rfloor+1}.$ Both $G_{3}$ and $G_{4}$ are $\{K_{r},P_{k}\}$-free and the lower bound follows.
For the upper bound, assume that $G$ is a $\{K_{r},P_{k}\}$-free graph with $\lfloor k/2\rfloor+1\leq r<k\leq n$. Let $u$ be a new vertex. The graph resulted from connecting $u$ to each vertex of $G$ is denoted by $G^{\prime}$. Notice that $G^{\prime}$ contains $n^{\prime}=n+1$ vertices and $e(G)+n$ edges. Additionally, $G^{\prime}$ is $\{K_{r+1},{\cal C}_{\geq k+1}\}$-free. Let $r^{\prime}=r+1$ and $k^{\prime}=k+1$. Observe that $r^{\prime}\geq\lfloor(k^{\prime}-1)/2\rfloor+2$, $r^{\prime}<k^{\prime}\leq n^{\prime}$, and $n^{\prime}-1=p(k^{\prime}-2)+q$.
If $k$ is odd and $k=3$, then $G$ is a matching as $G$ is $P_{3}$-free which is trivial. If $k$ is odd and $k\geq 5$, i.e., $k^{\prime}\geq 6$ is even, then Theorem 4 yields that
$e(G^{\prime})\leq p{\rm ex}(k^{\prime}-1,K_{r^{\prime}})+{\rm ex}(q+1,K_{r^{\prime}}).$
As the assumption of $k$ and $r$, to get $T(k^{\prime}-1,r^{\prime}-1)$, one can add a new vertex to $T(k-1,r-1)$ which is adjacent to all vertices in $T(k-1,r-1)$. It follows that ${\rm ex}(k^{\prime}-1,K_{r^{\prime}})={\rm ex}(k-1,K_{r})+k-1.$ Similarly, ${\rm ex}(q+1,K_{r^{\prime}})={\rm ex}(q,K_{r})+q.$ Recall $n=p(k-1)+q$. Thus
$e(G)=e(G^{\prime})-n\leq p{\rm ex}(k^{\prime}-1,K_{r^{\prime}})+{\rm ex}(q+1,K_{r^{\prime}})-n=p{\rm ex}(k-1,K_{r})+{\rm ex}(q,K_{r}).$
The case where $k$ is even can be proved similarly. $\square$
3.2 Proof of Theorem 4
We sketch the proof of Theorem 4 first. Let $G$ be a $\{K_{r},{\cal C}_{\geq k}\}$-free graph. We apply the induction on the number of blocks in $G$. For even $k$, as we can show $\max\{g_{r}(n,(k-2)/2,k),g_{r}(n,2,k)\}\leq f(n,k,r)$ (see Lemma 13 and Lemma 14), then the theorem follows from the subadditivity of $f(n,k,r)$ (see Lemma 11). For odd $k$, as $g_{r}(n,(k-1)/2,k)>f(n,k,r)$ for large $n$, then the case where there is a block with $g_{r}(b,(k-1)/2,k)$ edges will be handled separately. The other case in which each block $B$ contains $g_{r}(b,2,k)$ edges also follows from subadditivity of $f(n,k,r)$.
We begin to prove a number of technical lemmas. Readers can skip proofs of lemmas first and go the proof of Theorem 4 directly.
The next lemma gives a formula for the number of edges in a certain Turán graph.
Lemma 10
If $r\geq\lfloor(k-1)/2\rfloor+2$ and $r-1\leq n\leq k-1$, then ${\rm ex}(n,K_{r})=\binom{n}{2}-(n-r+1)$.
Proof: Notice that ${\rm ex}(n,K_{r})$ is the number of edges in the Turán graph $T(n,r-1)$. As the assumption $r\geq\lfloor(k-1)/2\rfloor+2\geq k/2+1$ and $r\leq n\leq k-1$ , each partite of $T(n,r-1)$ contains at most two vertices and there are $n-r+1$ partites with exactly two vertices. Thus
${\rm ex}(n,K_{r})=\binom{n}{2}-(n-r+1),$
as required. $\square$
Recall the following definition. For integers $n$ and $k$, if $n-1=p(k-2)+q$ with $q\leq k-3$, then
$f(n,k,r)=p{\rm ex}(k-1,K_{r})+{\rm ex}(q+1,K_{r}).$
We next prove the subadditivity of the function $f(n,k,r)$.
Lemma 11
Let $n-1=p(k-2)+q$, $n_{1}-1=p_{1}(k-2)+q_{1}$ and $n_{2}-1=p_{2}(k-2)+q_{2}$, where $n_{1}+n_{2}-1=n$ and $30\leq q,q_{1},q_{2}\leq k-$ 3. If $r\geq\lfloor(k-1)/2\rfloor+2$, then we have
$f(n_{1},k,r)+f(n_{2},k,r)\leq f(n,k,r).$
Proof: Let $G$ be the graph which consists of $p$ copies of $T(k-1,r-1)$ and one $T(q+1,r-1)$ all sharing a common vertex. Similarly, for $i=1,2$, let $G_{i}$ be the graph consists of $p_{i}$ copies of $T(k-1,r-1)$ and one $T(q_{i}+1,r-1)$ sharing a common vertex. By the definition, we have $f(n,k,r)=e(G)$ and $f(n_{i},k,r)=e(G_{i})$ for $i\in\{1,2\}$. As the assumption for $q_{1}$ and $q_{2}$, it follows that either $p_{1}+p_{2}=p$ or $p_{1}+p_{2}=p+1$.
If $p=p_{1}+p_{2}$, then $q=q_{1}+q_{2}$ in this case. Removing edges contained in all $T(k-1,r-1)$ from $G$, $G_{1}$, and $G_{2}$, we can see that it is left to show $t(q+1,r-1)\geq t(q_{1}+1,r-1)+t(q_{2}+1,r-1)$ with the assumption $q=q_{1}+q_{2}.$ As the assumption $r\geq\lfloor(k-1)/2\rfloor+2$ and $q_{i}\leq k-3$ for $i\in\{1,2\}$, there is at least one part from $T(q_{1}+1,r-1)$ and $T(q_{2}+1,r-1)$ with one vertex. If one identify such a vertex from $T(q_{1}+1,r-1)$ with the one from $T(q_{2}+1,r-1)$, then the resulting graph is $K_{r}$-free and with $q_{1}+q_{2}+1$ vertices and $t(q_{1}+1,r-1)+t(q_{2}+1,r-1)$ edges. Now the inequality $t(q+1,r-1)\geq t(q_{1}+1,r-1)+t(q_{2}+1,r-1)$ follows from the definition of the Turán graph.
If $p=p_{1}+p_{2}+1$, then $q_{1}+q_{2}=q+k-2$. It remains to establish $t(q_{1}+1,r-1)+t(q_{2}+1,r-1)\leq t(k-1,r-1)+t(q+1,r-1)$ under the condition $q_{1}+q_{2}=q+k-2$. Without loss of generality, we assume $q_{1}\geq q_{2}$. For $22\leq a\leq b\leq k-$ 2, we next show the fact that
$t(a,r-1)+t(b,r-1)\leq t(a-1,r-1)+t(b+1,r-1).$
As the assumption $r\geq\lfloor(t-1)/2\rfloor+2$, we observe that each partite in the Turán graph $T(x,r-1)$ for $12\leq x\leq k-$ 1 contains either two vertices or one vertex. To see the fact, starting from the vertex disjoint Turán graphs $T(a,r-1)$ and $T(b,r-1)$, we move one vertex $u$ from $T(a,r-1)$ to $T(b,r-1)$ such that the resulting graphs are $T(a-1,r-1)$ and $T(b+1,r-1)$. By the observation above, through this process, we remove at most $a-1$ edges and create at least $b-1$ edges (the vertex $u$ is contained in a partite with two vertices in $T(b+1,r-1)$). The fact follows easily as $a\leq b$. We can apply this fact recursively staring with $a=q_{1}+1$, $b=q_{2}+1$, and stop once one Turán graph contains exactly $q+1$ vertices. $\square$
Recall $g_{r}(n,a,k)=a(n-k+a)+{\rm ex}(k-a,K_{r})$. We first prove $g_{r}(n,2,k)\leq f(n,k,r)$. The following fact gives us the range in which $g_{r}(n_{,}2,k)>g_{r}(n,\lfloor(k-1)/2\rfloor,k)$.
Lemma 12
Assume that $n\geq k\geq 7$ and $\lfloor(k-1)/2\rfloor+2\leq r<k$. If $g_{r}(n_{,}2,k)>g_{r}(n,\lfloor(k-1)/2\rfloor,k)$, then $n\leq\tfrac{5k}{4}-1$.
Proof: Suppose that $n>\tfrac{5k}{4}-1$. We next show
$K=g_{r}(n,\lfloor(k-1)/2\rfloor,k)-g_{r}(n,2,k)\geq 0,$
which is a contradiction to the assumption. Recall that
$g_{r}(n,a,k)=(n-k+a)a+{\rm ex}(k-a,K_{r}).$
Moreover, Lemma 10 yields that ${\rm ex}(k-2,K_{r})=\binom{k-2}{2}-(k-r-1).$ Thus we have
$g_{r}(n,2,k)$ $=2(n-k+2)+{\rm ex}(k-2,K_{r})$
$=2n+\frac{k^{2}}{2}-\frac{11}{2}k+r+8.$
If $k$ is even, then we have $r\geq k/2+1$. It implies that
$g_{r}\left(n,\frac{k-2}{2},k\right)$ $=\frac{k-2}{2}\left(n-\frac{k+2}{2}\right)+{\rm ex}\left(\frac{k+2}{2},K_{r}\right)$
$\geq\frac{k-2}{2}\left(n-\frac{k+2}{2}\right)+\binom{k/2+1}{2}-1$
$=\frac{k-2}{2}n-\frac{k^{2}}{8}+\frac{k}{4}.$
Recall assumptions $n>\tfrac{5k}{4}-1$, $k\geq 8$ ($k$ is even and $k\geq 7$), and $r<k$. Thus
$K$ $\geq\frac{k-2}{2}n-\frac{k^{2}}{8}+\frac{k}{4}-\left(2n+\frac{k^{2}}{2}-\frac{11}{2}k+r+8\right)$
$=\left(\frac{k}{2}-3\right)n-\frac{5}{8}k^{2}+\frac{23}{4}k-r-8$
$>\left(\frac{k}{2}-3\right)\left(\frac{5}{4}k-1\right)-\frac{5}{8}k^{2}+\frac{23}{4}k-r-8$
$=\frac{3k}{2}-r-5$
$\geq\frac{3k}{2}-(k-1)-5$
$=k/2-4\geq 0,$
as desired.
If $k$ is odd, then we have $r\geq\frac{k+3}{2}.$ It follows that
$g_{r}\left(n,\frac{k-1}{2},k\right)$ $=\frac{k-1}{2}\left(n-\frac{k+1}{2}\right)+{\rm ex}\left(\frac{k+1}{2},K_{r}\right)$
$=\frac{k-1}{2}\left(n-\frac{k+1}{2}\right)+\binom{(k+1)/2}{2}$
$=\frac{k-1}{2}n-\frac{k^{2}}{8}+\frac{1}{8}.$
Therefore,
$K$ $=\frac{k-1}{2}n-\frac{k^{2}}{8}+\frac{1}{8}-\left(2n+\frac{k^{2}}{2}-\frac{11}{2}k+r+8\right)$
$=\frac{k-5}{2}n-\frac{5}{8}k^{2}+\frac{11}{2}k-r-\frac{63}{8}$
$>\frac{k-5}{2}\left(\frac{5k}{4}-1\right)-\frac{5}{8}k^{2}+\frac{11}{2}k-r-\frac{63}{8}$
$=\frac{15k}{8}-\frac{43}{8}-r$
$\geq\frac{15k}{8}-\frac{43}{8}-(k-1)$
$=\frac{7}{8}(k-5)>0,$
as desired. The proof is complete. $\square$
We next show $g_{r}(n,2,k)\leq f(n,k,r)$.
Lemma 13
If $16\leq k\leq n\leq\tfrac{5k}{4}-$ 1, $r\geq\lfloor(k-1)/2\rfloor+2$, and $n-1=k-2+q$ with $q\leq k-3$, then
$g_{r}(n,2,k)\leq f(n,k,r).$
Proof: As $k\leq n\leq\tfrac{5k}{4}-1$, it follows that $q\leq\tfrac{k}{4}.$ Since $r\geq\lfloor(k-1)/2\rfloor+2\geq k/2+1,$ we have $r-1\geq q+1$ and ${\rm ex}(q+1,K_{r})=\binom{q+1}{2}.$ By Lemma 10, we have
${\rm ex}(k-1,K_{r})=\binom{k-1}{2}-(k-r).$
Similarly, ${\rm ex}(k-2,K_{r})=\binom{k-2}{2}-(k-r-1)$. Then
$M$ $=f(n,k,r)-g_{r}(n,2,k)$
$=\binom{k-1}{2}-(k-r)+\binom{q+1}{2}-2(n-k+2)-\binom{k-2}{2}+(k-r-1)$
$=k-3+\binom{q+1}{2}-2(n-k+2).$
If $q\geq 2$, then $M\geq k-2(n-k+2)\geq\tfrac{k}{2}-2\geq 0$ as $k\geq 6$. As $n\geq k$, it is left to consider the case where $q=1$ and $n=k$. Note that $M\geq 0$ in this case since the assumption $k\geq 6$. $\square$
For even $k$, the next lemma shows $g_{r}(n,(k-2)/2,k)\leq f(n,k,r)$.
Lemma 14
Let $n-1=p(k-2)+q$ with $q\leq k-3$. If $k$ is even, $n\geq k\geq 6$, and $k/2+1\leq r\leq k-1$, then $g_{r}(n,(k-2)/2,k)\leq f(n,k,r)$.
Proof: Note that $(k-2)/2=\lfloor(k-1)/2\rfloor$ as $k$ is even. Additionally,
$g_{r}(n,(k-2)/2,k)=\frac{k-2}{2}(n-(k+2)/2)+{\rm ex}((k+2)/2,K_{r}).$
Notice that ${\rm ex}(k-1,K_{r})$ is the number of edges in the Turán graph $T(k-1,r-1)$. As the assumption $k/2+1\leq r\leq k-1$, by Lemma 10, we have
${\rm ex}(k-1,K_{r})=\binom{k-1}{2}-(k-r).$
Similarly, ${\rm ex}(q+1,K_{r})=\binom{q+1}{2}-(q-r+2)$ for $q+1\geq r$ and ${\rm ex}(q+1,K_{r})=\binom{q+1}{2}$ for $q+1\leq r-1$. Recall
$f(n,k,r)=p{\rm ex}(k-1,K_{r})+{\rm ex}(q+1,K_{r}).$
Case 1: $r\geq k/2+2$. Apparently,
${\rm ex}((k+2)/2,K_{r})=\binom{(k+2)/2}{2}\textrm{ and }g_{r}(n,(k-2)/2,k)=\frac{(k-2)n}{2}-\frac{k^{2}}{8}+\frac{k}{4}+1.$
If $q+1\geq r$, then
$f(n,k,r)$ $=p\binom{k-1}{2}-p(k-r)+\binom{q+1}{2}-(q-r+2)$
$=\frac{n-1-q}{k-2}\binom{k-1}{2}+\binom{q+1}{2}-\frac{n-1-q}{k-2}(k-r)-(q-r+2)$
$=\frac{(k-1)n}{2}-\frac{(q+1)(k-1)}{2}+\frac{q(q+1)}{2}-\frac{n-1-q}{k-2}(k-r)-(q-r+2)$
$\geq\frac{(k-1)n}{2}-\frac{(q+1)(k-1)}{2}+\frac{q(q+1)}{2}-\frac{n-1-q}{k-2}(\frac{k}{2}-2)-(q-r+2)\hfil\textrm{ (as }r\geq k/2+2)$
$=\frac{(k-2)n}{2}+\frac{(q+1)^{2}}{2}-\frac{(q+1)(k+1)}{2}+\frac{n-1-q}{k-2}+r-1$
$\geq\frac{(k-2)n}{2}-\frac{k^{2}}{8}-\frac{k}{4}+1+\frac{n-1-q}{k-2}+r-1\hfil(\textrm{ as }q+1\geq r\geq k/2+2)$
$\geq\frac{(k-2)n}{2}-\frac{k^{2}}{8}+\frac{k}{4}+1$
$=g_{r}(n,(k-2)/2,k).$
For $q+1\leq r-1$, we get that
$f(n,k,r)$ $=p\binom{k-1}{2}-p(k-r)+\binom{q+1}{2}$
$=\frac{n-1-q}{k-2}\binom{k-1}{2}+\binom{q+1}{2}-\frac{n-1-q}{k-2}(k-r)$
$=\frac{(k-1)n}{2}-\frac{(q+1)(k-1)}{2}+\frac{q(q+1)}{2}-\frac{n-1-q}{k-2}(k-r)$
$\geq\frac{(k-1)n}{2}-\frac{(q+1)(k-1)}{2}+\frac{q(q+1)}{2}-\frac{n-1-q}{k-2}(\frac{k}{2}-2)\hfil\textrm{ (as }r\geq k/2+2)$
$=\frac{(k-2)n}{2}+\frac{(q+1)^{2}}{2}-\frac{(q+1)(k-1)}{2}+\frac{n-1-q}{k-2}$
$\geq\frac{(k-2)n}{2}-\frac{k^{2}}{8}+\frac{k}{4}+\frac{n-1-q}{k-2}\hfil(\textrm{ as the minimum is achievd for }q+1\in\{k/2,k/2-1\})$
$\geq\frac{(k-2)n}{2}-\frac{k^{2}}{8}+\frac{k}{4}+1\hfil(\textrm{ as }p=\frac{n-1-q}{k-2}\geq 1)$
$=g_{r}(n,(k-2)/2,k).$
Case 2: $r=k/2+1$. Observe that
${\rm ex}((k+2)/2,K_{r})=\binom{(k+2)/2}{2}-1\textrm{ and }g_{r}(n,k/2-1,k)=\frac{(k-2)n}{2}-\frac{k^{2}}{8}+\frac{k}{4}$
Repeating the argument in Case 1, we can show
$f(n,k,r)\geq g_{r}(n,k/2-1,k)$
similarly. $\square$
As we did before, let $t=\lfloor(k-1)/2\rfloor$. Recall
$g_{r}(n,a,k)=(n-k+a)a+{\rm ex}(k-a,K_{r})$
and
$f(n,k,r)=p{\rm ex}(k-1,K_{r})+{\rm ex}(q+1,K_{r}).$
Proof of Theorem 4: even $k$: To see the lower bound, observe that $F(n,k,r)$ is $\{K_{r},{\cal C}_{\geq k}\}$-free and $e(F(n,k,r))=f(n,k,r)$. The lower bound follows.
For the upper bound, let $G$ be an $n$-vertex $\{K_{r},{\cal C}_{\geq k}\}$-free graph with maximum number of edges. We claim that $G$ is connected. Otherwise, if $G$ is not connected, then the graph resulted from adding an edge to connect two connected components is still $\{K_{r},{\cal C}_{\geq k}\}$-free, but has more edges than $G$. This is a contradiction to the definition of $G$. Thus $G$ is connected. We apply the induction on the number of blocks contained in $G$ to show the upper bound. If $G$ contains one block ($G$ is 2-connected), then $e(G)\leq{\rm ex}(n,K_{r})=f(n,k,r)$ for $n\leq k-1$ and
$e(G)\leq\max\{g_{r}(n,2,k),g_{r}(n,t,k)\}\leq f(n,k,r)$
for $n\geq k$. The inequality above follows from the combination of Theorem 7, Lemma 12, Lemma 13, and Lemma 14.
Assume that $G$ contains at least two blocks. Let $B$ be an end block of $G$ and $b$ be the cut vertex. Let $G^{\prime}=(G-B)\cup\{b\}$ and $n^{\prime}=|V(G^{\prime})|$. Then $G^{\prime}$ contains fewer blocks than $G$. If $n^{\prime}<k$, then $e(G^{\prime})\leq{\rm ex}(n,K_{r})=f(n^{\prime},k,r)$. Otherwise, by induction hypothesis, $e(G^{\prime})\leq f(n^{\prime},k,r)$. Additionally, $e(B)\leq f(n-n^{\prime}+1,k,r)$ by using the proof for the base case. Therefore,
$e(G)=e(G^{\prime})+e(B)$ $\leq f(n^{\prime},k,r)+f(n-n^{\prime}+1,k,r)$
$\leq f(n,k,r).$
The proof of the even case is complete. $\square$
Proof of Theorem 4: odd $k$: For the lower bound, we already showed that $F(n,k,r)$ is $\{K_{r},{\cal C}_{\geq k}\}$-free. Note that the clique number of $G_{1}$ is $t+1$ which is less that $r$ by the assumption of $r$. In addition, a longest cycle contains at most $k-1$ edges. Thus $G_{1}$ is also $\{K_{r},{\cal C}_{\geq k}\}$-free. The lower bound is proved.
For the upper bound, we assume $k\geq 7$ for a moment. As we did above, let $G$ be an $n$-vertex $\{K_{r},{\cal C}_{\geq k}\}$-free graph with maximum number of edges. Then $G$ is connected. Since both $K_{r}$ and ${\cal C}_{\geq k}$ are 2-connected, we can further assume that the subgraph induced by each block contains the maximum number of edges. In other words, if $B$ is a block and $b=|V(B)|$, then either $e(G[B])={\rm ex}(b,K_{r})$ for $12\leq b\leq k-$ 1 or $e(G[B])=\max\{g_{r}(b,2,k),g_{r}(b,t,k)\}$ for $b\geq k$. Apparently, each block $B$ is ${\cal C}_{\geq k}$-free. Note that $(k-1)/2=\lfloor(k-1)/2\rfloor=t$ as $k$ is odd. Theorem 2 yields that $e(B)\leq t(|V(B)|-1)$ no matter the size of $B$.
Assume $B_{1},\ldots,B_{j}$ are blocks of $G$. For each $1\leq i\leq j$, let $b_{i}$ be the number of vertices in $B_{i}$. Consider a rooted block tree with $B_{1}$ as the root. Observe that each block $B_{i}$ brings exactly $b_{i}-1$ new vertices and then $\sum_{i=2}^{j}(b_{i}-1)=n-b_{1}$. For the case where $G$ contains a block, say $B_{1}$, such that $e(B_{1})=g_{r}(n,t,k)$, then
$e(G)=$ $\sum_{i=1}^{j}e(B_{i})=e(B_{1})+\sum_{i=2}^{j}e(B_{i})$
$\leq e(B_{1})+\sum_{i=2}^{j}t(b_{i}-1)$
$=g_{r}(b_{1},t,k)+\sum_{i=2}^{j}t(b_{i}-1)$
$=g_{r}(b_{1},t,k)+t(n-b_{1})$
$=\binom{t}{2}+t(b_{1}-1)-t(t-1)+t(n-b_{1})$
$=\binom{t}{2}+t(n-1)-t(t-1)$
$=g_{r}(n,t,k).$
It remains to consider the case where each block $B_{i}$ in $G$ satisfies $e(B_{i})\neq g_{r}(n,t,k)$. We apply the induction on the number of blocks in $G$ to show $e(G)\leq f(n,k,r)$. For the case where $G$ contains one block (i.e., $G$ is 2-connected), if $n\geq k$, then $e(G)\leq g_{r}(n,2,k)\leq f(n,k,r)$ by the assumption of $G$ and Lemma 13. If $n<k$, then $e(G)\leq{\rm ex}(n,K_{r})=f(n,k,r)$. For the case where $G$ contains at least two blocks, let $B$ be an end block of $G$ and $b$ be the cut vertex. Set $G^{\prime}=(G-B)\cup\{b\}$. Then $G^{\prime}$ contains fewer blocks than $G$. Let $n^{\prime}=|V(G^{\prime})|$. Then $e(G^{\prime})\leq f(n^{\prime},k,r)$ for $n^{\prime}\geq k$ by the induction hypothesis and $e(G^{\prime})\leq{\rm ex}(n,K_{r})=f(n^{\prime},k,r)$ for $n^{\prime}<k$. Moreover, repeating the proof for base case, we can show $e(G[B])\leq f(b,k,r)=f(n-n^{\prime}+1,k,r)$. By induction hypothesis and Lemma 11, it follows that
$e(G)=e(G^{\prime})+e(B)$ $\leq f(n^{\prime},k,r)+f(n-n^{\prime}+1,k,r)$
$\leq f(n,k,r).$
The upper bound is proved for $k\geq 7$. It remains to consider the case where $k=5$. In this case, note that $r=4$ and $e(B_{i})=g_{4}(b_{i},2,5)$ whenever $b_{i}\geq 5$. Moreover, $e(B_{i})\leq 2(b_{i}-1)$ still holds. We can repeat the argument for $k\geq 7$ to show the upper bound. The proof for Theorem 4 is complete. $\square$
3.3 Proof of Theorem 5
To prove Theorem 5, let $G$ be a $\{K_{r},{\cal C}_{\geq k}\}$-free graph. We also apply the induction on the number blocks contained in $G$. Since Theorem 5 only works for large $n$, we need to establish an upper bound for the number of edges in small blocks. The proof will be split into two cases depending on whether there is a large block.
We again fix $t=\lfloor(k-1)/2\rfloor$. To prove the upper bound in Theorem 5, we need two more specific lemmas.
Lemma 15
If $2\leq n<k,k\geq 5$ and $13\leq r\leq t+$ 1, then ${\rm ex}(n,K_{r})\leq(t-1/k)(n-1).$
Proof: For the case that $n\leq k-3$, we have
${\rm ex}(n,K_{r})\leq\binom{n}{2}=\frac{n(n-1)}{2}\leq\frac{(k-3)(n-1)}{2}\leq(t-\frac{1}{k})(n-1)$
as $t=\lfloor(k-1)/2\rfloor.$ It remains to show that the inequality holds for $n=\{k-2,k-1\}.$ If $n=k-2$ and $k$ is odd, then we also have ${\rm ex}(n,K_{r})\leq(t-\frac{1}{k})(n-1)$ by the argument above. If $n=k-2$ and $k$ is even, then $r-1\leq t=(k-2)/2$ and each partite of $T(k-2,r-1)$ contains at least two vertices. As $r-1\geq 2$ and $11-3/k<$ 1, we have
${\rm ex}(k-2,K_{r})\leq\frac{(k-2)(k-3)}{2}-(r-1)\leq\frac{k-2}{2}(k-3)-(1-3/k)=\left(t-\frac{1}{k}\right)(k-3).$
For the case that $n=k-1$, as $r-1\leq t$, we have
${\rm ex}(k-1,K_{r})\leq\left(1-\frac{1}{r-1}\right)\frac{(k-1)^{2}}{2}\leq\left(1-\frac{1}{t}\right)\frac{(k-1)^{2}}{2}\leq\left(t-\frac{1}{k}\right)(k-2).$
The last inequality can be verified directly by noting that $t=\lfloor(k-1)/2\rfloor$ and $k\geq 5.$square
Lemma 16
If $5\leq k\leq n$ and $13\leq r\leq t+$ 1, then
${\rm ex}(k-1,K_{r})+\left(t-\frac{1}{2}\right)(n-k+1)\leq\left(t-\frac{1}{k}\right)(n-1).$
Proof: Lemma 15 gives that ${\rm ex}(k-1,K_{r})\leq(t-\tfrac{1}{k})(k-2)$. Apparently, $(t-\tfrac{1}{2})(n-k+1)\leq(t-\tfrac{1}{k})(n-k+1)$ as $k\geq 5$. The lemma follows. $\square$
We are now ready to prove Theorem 5.
Proof of Theorem 5: For the lower bound, as the clique number of $G_{2}$ is $r-1$ and a longest cycle contains at most $12t\leq k-$ 1 edges. Thus $G_{2}$ is $\{K_{r},{\cal C}_{\geq k}\}$-free and the lower bound follows.
To make the notation simple, we define
$g^{\prime}_{r}(n,t,k)=(n-t)t+{\rm ex}(t,K_{r-1}).$
Note that
$g^{\prime}_{r}(n,t,k)=t(n-1)-t(t-1)+{\rm ex}(t,K_{r-1})\leq t(n-1).$
We next prove the upper bound. Let $G$ be an $n$-vertex $\{K_{r},{\cal C}_{\geq k}\}$-free graph with maximum number of edges, where $n\geq\tfrac{k^{3}}{4}$. As we did in the proof of Theorem 4, we can assume that $G$ is connected. Let $B_{1},\ldots,B_{j}$ be blocks of $G$. For $1\leq i\leq j$, let $b_{i}$ be the size of $B_{i}$.
If $b_{i}\geq\tfrac{k^{2}}{2}$, then Theorem 8 gives us $e(B_{i})\leq t(b_{i}-t)+{\rm ex}(t,K_{r-1})=g^{\prime}_{r}(b_{i},t,k).$ If $k\leq b_{i}<\tfrac{k^{2}}{2}$, then by Lemma 9, either $e(B_{i})\leq t(b_{i}-t)+{\rm ex}(t,K_{r-1})=g^{\prime}_{r}(b_{i},t,k)$ or
$e(B_{i})\leq\max\{(t-1/2)(b_{i}-1),{\rm ex}(k-1,K_{r})+(t-\tfrac{1}{2})(b_{i}-k+1)\}\leq(t-1/k)(b_{i}-1).$
The last inequality follows from Lemma 16. If $b_{i}<k$, then we also have
$e(B_{i})\leq{\rm ex}(b_{i},K_{r})\leq(t-1/k)(b_{i}-1).$
by Lemma 15. Note that $e(B_{i})\leq t(b_{i}-1)$ holds for any $1\leq i\leq j$ as $B_{i}$ is ${\cal C}_{\geq k}$-free and Theorem 2.
For the case where there is a $B_{i}$, say $B_{1}$, such that $b_{1}\geq\tfrac{k^{2}}{2}$, then $e(B_{1})\leq g^{\prime}_{r}(b_{1},t,k)$. Note that $\sum_{i=2}^{j}(b_{i}-1)=n-b_{1}$ by considering a rooted block tree with $B_{1}$ as the root. We can upper bound $e(G)$ as follows:
$e(G)$ $=\sum_{i=1}^{j}e(B_{i})$
$=e(B_{1})+\sum_{i=2}^{j}e(B_{i})$
$\leq e(B_{1})+\sum_{i=2}^{j}t(b_{i}-1)$
$=g_{r}^{\prime}(b_{1},t,k)+\sum_{i=2}^{j}t(b_{i}-1)$
$=g_{r}^{\prime}(b_{1},t,k)+t(n-b_{1})$
$=t(b_{1}-1)-t(t-1)+{\rm ex}(t,K_{r-1})+t(n-b_{1})$
$=t(n-1)-t(t-1)+{\rm ex}(t,K_{r-1})$
$=t(n-t)+{\rm ex}(t,K_{r-1}).$
It remains to consider the case where $b_{i}<\tfrac{k^{2}}{2}$ for each $1\leq i\leq j$. If there is a $B_{i}$ such that $k\leq b_{i}\leq\tfrac{k^{2}}{2}$ and $e(B_{i})\leq g^{\prime}_{r}(b_{i},t,k)$, then we can repeat the argument above to show the desired upper bound. We are left to show the upper bound for the case where $b_{i}<\tfrac{k^{2}}{2}$ and $e(B_{i})\leq(t-1/k)(b_{i}-1)$ for each $1\leq i\leq j$. For this case, we have
$e(G)$ $=\sum_{i=1}^{j}e(B_{i})$
$\leq\sum_{i=1}^{j}(t-1/k)(b_{i}-1)$
$=(t-1/k)(n-1)$
$\leq t(n-1)-t(t-1)+{\rm ex}(t,K_{r-1})$
$=t(n-t)+{\rm ex}(t,K_{r-1}),$
the second-to-last inequality follows from the assumption $n\geq\tfrac{k^{3}}{4}$. The proof is complete. $\square$
4 Concluding remarks
In this paper, we proved results on the Turán number of =${\cal H}=\{K_{r},{\cal C}_{\geq k}\}$. For $r\geq\lfloor(k-1)/2\rfloor+2$, we obtained the value of ${\rm ex}(n,{\cal H})$ for all $n$. However, for $13\leq r\leq\lfloor(k-1)/2\rfloor+$ 1, we were only able to show the value of ${\rm ex}(n,{\cal H})$ for large $n$. A natural question is to determine ${\rm ex}(n,{\cal H})$ for small $n$ in this case. Let us recall lower bound constructions in this paper (Theorem 4 and Theorem 5). For the graph $F(n,k,r)$, we start with a ${\cal C}_{\geq k}$-free graph with the maximum number of edges and turn it $K_{r}$-free. To construct $G_{2}$, we modify the Turán graph $T(n,r-1)$ so that the sum of size of smallest $(r-2)$-partite is at most $\lfloor(k-1)/2\rfloor$ (to ensure it is ${\cal C}_{\geq k}$-free). The answer to the question above requires new ideas to construct ${\cal H}$-free graphs.
Note that upper bounds for ${\rm ex}(n,P_{k})$ and ${\rm ex}(n,{\cal C}_{\geq k})$ in Theorem 1 and Theorem 2 are only tight for particular $n$. However, these upper bounds are widely applicable because they are very simple. Therefore, it is desired to show such kind upper bound for ${\rm ex}(n,\{K_{r},{\cal C}_{\geq k}\})$. Unfortunately, this kind of upper bound does not exist in general as Theorem 4 involves two extremal graphs for odd $k$. Therefore, in comparison to the Turán function ${\rm ex}(n,{\cal C}_{\geq k})$, the Turán function ${\rm ex}(n,\{K_{r},C_{\geq k}\})$ indeed behaves differently.
As we mentioned before, the result on the Turán number of $P_{k}$ is a simple corollary of the one for ${\cal C}_{\geq k}$, one may ask whether it is still the case for $K_{r}$-free graphs. The answer is positive for $r\geq\lfloor(k-1)/2\rfloor+2$ as Theorem 6. For $r\leq\lfloor(k-1)/2\rfloor+1$, we are not able to prove such kind result. One possible reason is that ${\rm ex}(n,\{K_{r},{\cal C}_{\geq k}\})$ is not known for small $n$. If one can determine ${\rm ex}(n,\{K_{r},{\cal C}_{\geq k}\})$ for all $n$, then it is very likely to provide a positive answer to the question mentioned above.
It is obvious that ${\rm ex}_{2\textrm{-conn}}(n,{\cal H})\leq{\rm ex}(n,{\cal H})$ for any ${\cal H}$ by definitions. For =${\cal H}=\{K_{r},{\cal C}_{\geq k}\}$, if $r=k/2+1$ for even $k$ and $n-1=p(k-2)+q$ with $p\geq 1$ and $q\in\{k/2-1,k/2-2\}$, then $f(n,k,r)=g_{r}(n,k/2-1,k)$ and $\max\{g_{r}(n,2,k),g_{r}(n,k/2-1,k)\}=g_{r}(n,k/2-1,k)$ by the direct computation. Therefore, ${\rm ex}_{2\textrm{-conn}}(n,{\cal H})={\rm ex}(n,{\cal H})$ by Theorem 4 and Theorem 7, which indicates that $r=\lfloor(k-1)/2\rfloor+2$ is indeed a critical case.